One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
I have tried this below, but I am unsure if I have done this correctly. This is so confusing, and driving me crazy as I try to figure this out.
x(x+12)=x+2x

Answer by stanbon(48568)

One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
-------------
smaller number = x
larger number = x+12
EQUATIOn:
x + 2(x+12) = 39
x+2x+24 = 39
3x = 15
x = 5 (smaller number)
x+12 = 17 (larger number)
===========
Cheers,
Stan H.